# On the Mode-Seeking Properties of Langevin Dynamics

### v1

@@ 1 Introduction
A central task in unsupervised learning involves learning the underlying probability distribution of training data and efficiently generating new samples from the distribution. Score-based generative modeling (SGM) [1] has achieved state-of-the-art performance in various learning tasks including image generation [2, 3, 4, 5, 6, 7], audio synthesis [8, 9], and video generation [10, 11]. In addition to the successful empirical results, the convergence analysis of SGM has attracted significant attention in the recent literature [12, 13, 14, 15, 16].
@@
Stochastic gradient Langevin dynamics (SGLD) [17], as a fundamental methodology to implement and interpret SGM, can produce samples from the (Stein) score function of a probability density, i.e., the gradient of the log probability density function with respect to data. It has been widely recognized that a pitfall of SGLD is its slow mixing rate [18, 19, 20]. Specifically, [2] shows that under a multi-modal data distribution, the samples from Langevin dynamics may have an incorrect relative density across the modes. Based on this finding, [2] proposes anneal Langevin dynamics, which injects different levels of Gaussian noise into the data distribution and samples with SGLD on the perturbed distribution. While outputting the correct relative density across modes can be challenging for SGLD, a natural question is whether SGLD would be able to find all the modes of a multi-modal distribution.
@@
In this work, we study this question by analyzing the mode-seeking properties of SGLD. The notion of mode-seekingness [21, 22, 23, 24] refers to the property that a generative model captures only a subset of the modes of a multi-modal distribution. We note that a similar problem, known as metastability, has been studied in the context of Langevin diffusion, a continuous-time version of SGLD described by stochastic differential equation (SDE) [25, 26, 27]. Specifically, [25] gave a sharp bound on the mean hitting time of Langevin diffusion and proved that it may require exponential (in the space dimensionality $d$) time for transition between modes. Regarding discrete SGLD, [20] constructed a probability distribution whose density is close to a mixture of two well-separated isotropic Gaussians, and proved that SGLD could not find one of the two modes within an exponential number of steps. However, further exploration of mode-seeking tendencies of SGLD and its variants such as annealed Langevin dynamics for general distributions is still lacking in the literature.
@@
In this work, we theoretically formulate and demonstrate the potential mode-seeking tendency of SGLD. We begin by analyzing the convergence under a variety of Gaussian mixture probability distributions, under which SGLD could fail to visit all the mixture components within sub-exponential steps (in the data dimension). Subsequently, we generalize this result to mixture distributions with sub-Gaussian modes. This generalization extends our earlier result on Gaussian mixtures to a significantly larger family of mixture models, as the sub-Gaussian family includes any distribution over an $\ell_{2}$-norm-bounded support set. Furthermore, we extend our theoretical results to anneal Langevin dynamics with bounded noise scales.
@@
To reduce SGLD’s large iteration complexity shown under a high-dimensional input vector, we propose Chained Langevin Dynamics (Chained-LD). Since SGLD could suffer from the curse of dimensionality, we decompose the sample ${\mathbf{x}}\in{\mathbb{R}}^{d}$ into $d/Q$ patches ${\mathbf{x}}^{(1)},\cdots,{\mathbf{x}}^{(d/Q)}$, each of constant size $Q$, and sequentially generate every patch ${\mathbf{x}}^{(q)}$ for all $q\in[d/Q]$ statistically conditioned on previous patches, i.e., $P({\mathbf{x}}^{(q)}\mid{\mathbf{x}}^{(0)},\cdots{\mathbf{x}}^{(q-1)})$. The combination of all patches generated from the conditional distribution faithfully follows the probability density $P({\mathbf{x}})$, while learning each patch requires less cost due to the reduced dimension. We also provide a theoretical analysis of Chained-LD by reducing the convergence of a $d$-dimensional sample to the convergence of each patch.
@@
Finally, we present the results of several numerical experiments to validate our theoretical findings. For synthetic experiments, we consider moderately high-dimensional Gaussian mixture models, where the vanilla and annealed Langevin dynamics could not find all the components within a million steps, while Chained-LD could capture all the components with correct frequencies in ${\mathcal{O}}(10^{4})$ steps. For experiments on real image datasets, we consider a mixture of two modes by using the original images from MNIST/Fashion-MNIST training dataset (black background and white digits/objects) as the first mode and constructing the second mode by i.i.d. flipping the images (white background and black digits/objects) with probability 0.5. Following from [2], we trained a Noise Conditional Score Network (NCSN) to estimate the score function. Our numerical results indicate that vanilla Langevin dynamics can fail to capture the two modes, as also observed by [2]. On the other hand, Chained-LD was capable of finding both modes regardless of initialization. We summarize the contributions of this work as follows:
@@
  • Theoretically studying the mode-seeking properties of vanilla and annealed Langevin dynamics,
@@
  • Numerically comparing the mode-seeking properties of vanilla, annealed, and chained Langevin dynamics.
@@ 2 Related Works
Langevin Dynamics: The convergence guarantees for Langevin diffusion, a continuous version of Langevin dynamics, are classical results extensively studied in the literature [28, 29, 30, 31]. Langevin dynamics, also known as Langevin Monte Carlo, is a discretization of Langevin diffusion typically modeled as a Markov Chain Monte Carlo [17]. For unimodal distributions, e.g., the probability density function that is log-concave or satisfies log-Sobolev inequality, the convergence of Langevin dynamics is provably fast [32, 33, 34]. However, for multimodal distributions, the non-asymptotic convergence analysis is much more challenging [35]. [19] gave an upper bound on the convergence time of Langevin dynamics for arbitrary non-log-concave distributions with certain regularity assumptions, which, however, could be exponentially large without imposing more restrictive assumptions. [20] studied the special case of a mixture of Gaussians of equal variance and provided heuristic analysis of sampling from general non-log-concave distributions.
@@
Score-based Generative Modeling: Since [40] proposed sliced score matching which can train deep models to learn the score functions of implicit probability distributions on high-dimensional data, score-based generative modeling (SGM) has been going through a spurt of growth. Annealed Langevin dynamics [2] estimates the noise score of the probability density perturbed by Gaussian noise and utilizes stochastic gradient Langevin dynamics to generate samples from a sequence of decreasing noise scales. [3] conducted a heuristic analysis of the effect of noise levels on the performance of annealed Langevin dynamics. Denoising diffusion probabilistic model (DDPM) [4] incorporates a step-by-step introduction of random noise into data, followed by learning to reverse this diffusion process in order to generate desired data samples from the noise. [1] unified anneal Langevin dynamics and DDPM via a stochastic differential equation. A recent line of work focuses on the non-asymptotic convergence guarantees for SGM with an imperfect score estimation under various assumptions on the data distribution [41, 42, 12, 14, 43, 15, 16].
@@ 3.1 Langevin Dynamics
${\,\mathrm{d}}{\mathbf{x}}_{t}=-\nabla_{\mathbf{x}}\log P({\mathbf{x}}_{t}){\,\mathrm{d}}t+\sqrt{2}{\,\mathrm{d}}{\mathbf{w}}_{t},$
where ${\mathbf{w}}_{t}$ is the Wiener process on ${\mathbb{R}}^{d}$. To generate samples from Langevin diffusion, [17] proposed stochastic gradient Langevin dynamics (SGLD), a discretization of the SDE for $T$ iterations. Each iteration of SGLD is defined as
@@
and running SGLD on the perturbed data distribution $P_{{\sigma}_{t}}({\mathbf{x}})$ with gradually decreasing noise levels $\left\{{{\sigma}_{t}}\right\}_{t\in[T]}$, i.e.,
@@
where $\delta_{t}$ is the step size and $\bm{\epsilon}_{t}\sim{\mathcal{N}}(\bm{0}_{d},{\bm{I}}_{d})$ is Gaussian noise. When the noise level ${\sigma}$ is vanishingly small, the perturbed distribution is close to the true distribution, i.e., $P_{\sigma}({\mathbf{x}})\approx P({\mathbf{x}})$. Since we do not have direct access to the (perturbed) score function, [2] proposed the Noise Conditional Score Network (NCSN) ${\mathbf{s}}_{\bm{\theta}}({\mathbf{x}},{\sigma})$ to jointly estimate the scores of all perturbed data distributions, i.e.,
Assuming the NCSN has enough capacity, ${\mathbf{s}}_{\bm{\theta}^{*}}({\mathbf{x}},{\sigma})$ minimizes the loss ${\mathcal{L}}\left(\bm{\theta};\left\{{{\sigma}_{t}}\right\}_{t\in[T]}\right)$ if and only if ${\mathbf{s}}_{\bm{\theta}^{*}}({\mathbf{x}},{\sigma}_{t})=\nabla_{\mathbf{x}}\log P_{{\sigma}_{t}}({\mathbf{x}})$ almost surely for all $t\in[T]$.
@@ 3.2 Multi-Modal Distributions
We remark that sub-Gaussian distributions include a wide variety of distributions such as Gaussian distributions and any distribution within a bounded $\ell_{2}$-norm distance from the mean ${\bm{\mu}}$. From equation 2 we note that the perturbed distribution is the convolution of the original distribution and a Gaussian random variable, i.e., for random variables ${\mathbf{z}}\sim p$ and ${\mathbf{t}}\sim{\mathcal{N}}(\bm{0}_{d},{\bm{I}}_{d})$, their sum ${\mathbf{z}}+{\mathbf{t}}\sim p_{\sigma}$ follows the perturbed distribution with noise level ${\sigma}$. Therefore, a perturbed (sub)Gaussian distribution remains (sub)Gaussian. We formalize this property in Proposition 1 and defer the proof to Appendix A for completeness.
Proposition 1.
@@ 4 Theoretical Analysis of the Mode-Seeking Properties of Langevin Dynamics
In this section, we theoretically investigate the mode-seeking properties of vanilla and annealed Langevin dynamics. We begin with analyzing Langevin dynamics in Gaussian mixtures.
@@ 4.1 Langevin Dynamics in Gaussian Mixtures
Regarding the first requirement ${\nu}_{i}<{\nu}_{0}$, we first note that the probability density $p({\mathbf{z}})$ of a Gaussian distribution ${\mathcal{N}}({\bm{\mu}},{\nu}^{2}{\bm{I}}_{d})$ decays exponentially in terms of $\frac{\left\|{{\mathbf{z}}-{\bm{\mu}}}\right\|^{2}}{{\nu}^{2}}$. When a state ${\mathbf{z}}$ is sufficiently far from all modes (i.e., $\left\|{{\mathbf{z}}}\right\|\gg\left\|{{\bm{\mu}}_{i}}\right\|$), the Gaussian distribution with the largest variance (i.e., $P^{(0)}$ in Assumption 1) dominates all other modes because $\frac{\left\|{{\mathbf{z}}-{\bm{\mu}}_{0}}\right\|^{2}}{{\nu}_{0}^{2}}\approx\frac{\left\|{{\mathbf{z}}}\right\|^{2}}{{\nu}_{0}^{2}}\gg\frac{\left\|{{\mathbf{z}}}\right\|^{2}}{{\nu}_{i}^{2}}\approx\frac{\left\|{{\mathbf{z}}-{\bm{\mu}}_{i}}\right\|^{2}}{{\nu}_{i}^{2}}$. We call such mode $P^{(0)}$ the universal mode. Therefore, if ${\mathbf{z}}$ is initialized far from all modes, it can only converge to the universal mode because the gradient information of other modes is masked. Once ${\mathbf{z}}$ enters the universal mode $P^{(0)}$, if the step size $\delta_{t}$ of Langevin dynamics is small (i.e., $\delta_{t}\leq{\nu}_{0}^{2}$), it would take exponential steps to escape the local mode $P^{(0)}$; while if the step size is large (i.e., $\delta_{t}>{\nu}_{0}^{2}$), the state ${\mathbf{z}}$ would again be far from all modes and thus the universal mode $P^{(0)}$ dominates all other modes. Hence, ${\mathbf{z}}$ can only visit the universal mode unless the stochastic noise $\bm{\epsilon}_{t}$ miraculously leads it to the region of another mode. In addition, it can be verified that $\log\left(\frac{{\nu}_{i}^{2}}{{\nu}_{0}^{2}}\right)-\frac{{\nu}_{i}^{2}}{2{\nu}_{0}^{2}}+\frac{{\nu}_{0}^{2}}{2{\nu}_{i}^{2}}$ is a positive constant for ${\nu}_{i}<{\nu}_{0}$, thus the second requirement of Assumption 1 essentially represents $\left\|{{\bm{\mu}}_{i}-{\bm{\mu}}_{0}}\right\|^{2}\leq{\mathcal{O}}(d)$. We formalize the intuition in Theorem 1 and defer the proof to Appendix A.1.
@@
We note that $\left\|{{\mathbf{x}}_{t}-{\bm{\mu}}_{i}}\right\|^{2}>\frac{{\nu}_{0}^{2}+{\nu}_{\max}^{2}}{2}d$ is a strong notion of mode-seekingness, since the probability density of mode $P^{(i)}={\mathcal{N}}({\bm{\mu}}_{i},{\nu}_{i}^{2}{\bm{I}}_{d})$ concentrates around the $\ell_{2}$-norm ball $\left\{{{\mathbf{z}}:\left\|{{\mathbf{z}}-{\bm{\mu}}_{i}}\right\|^{2}\leq{\nu}_{i}^{2}d}\right\}$. This notion can also easily be translated into a lower bound in terms of other distance measures such as total variation distance and Wasserstein 2-distance. Moreover, in Theorem 2 we extend the result to annealed Langevin dynamics with bounded noise level, and the proof is deferred to Appendix A.2.
Theorem 2.

### v2

@@ 1 Introduction
Langevin dynamics, a.k.a. Langevin Monte Carlo, is a well-known Markov Chain Monte Carlo (MCMC) sampling methodology that has been widely used to implement and interpret score-based generative modeling. It can produce samples from the (Stein) score function of a probability density, i.e., the gradient of the log probability density function with respect to data. It has been widely recognized that a pitfall of Langevin dynamics is its slow mixing rate [1, 2, 3]. Specifically, [4] shows that under a multi-modal data distribution, the samples from Langevin dynamics may have an incorrect relative density across the modes. Based on this finding, [4] proposes anneal Langevin dynamics, which injects different levels of Gaussian noise into the data distribution and samples with Langevin dynamics on the perturbed distribution. While outputting the correct relative density across modes can be challenging for Langevin dynamics, a natural question is whether Langevin dynamics would be able to find all the modes of a multi-modal distribution.
@@
In this work, we study this question by analyzing the mode-seeking properties of Langevin dynamics. The notion of mode-seekingness [5, 6, 7, 8] refers to the property that a generative model captures only a subset of the modes of a multi-modal distribution. We note that a similar problem, known as metastability, has been studied in the context of Langevin diffusion, a continuous-time version of Langevin dynamics described by stochastic differential equation (SDE) [9, 10, 11]. Specifically, [9] gave a sharp bound on the mean hitting time of Langevin diffusion and proved that it may require exponential (in the space dimensionality $d$) time for transition between modes. Regarding discrete Langevin dynamics, [3] constructed a probability distribution whose density is close to a mixture of two well-separated isotropic Gaussians, and proved that Langevin dynamics could not find one of the two modes within an exponential number of steps. However, further exploration of the mode-seeking tendencies of Langevin dynamics and its variants for general distributions is still lacking in the literature.
In this work, we study Langevin dynamics under multi-modal distributions in a slightly different setting from the standard theory literature on sampling. As illustrated in Figure 1, the existing theoretical literature commonly considers a mixture of well-separated modes with bounded variance. On the other hand, in our analysis, we consider a low-density high-variance mode (referred to as Mode 0 or $P^{(0)}$) surrounding the other modes and filling up the space between the modes. Note that Langevin dynamics relies on the score function (i.e., the gradient of log-pdf) to search for the modes. For a mixture model with no in-between modes (Figure 1.a), Langevin dynamics could carry information about the direction towards the closest mode, even if they are far from all the modes. However, assuming an in-between mode $P^{(0)}$ with high variance (Figure 1.b), the gradient information outside the support set of the low-variance modes will be dominated by $P^{(0)}$, despite the minor overall mass of $P^{(0)}$. As a result, one can expect that the dynamics would randomly explore a large volume in ${\mathbb{R}}^{d}$ until finding a low-variance mode, which can take a significant time.
@@
To theoretically formulate and demonstrate the potential mode-seeking tendency of Langevin dynamics, we begin by analyzing the convergence for mixture distributions of Gaussian modes, under which Langevin dynamics could fail to visit all the mixture components within sub-exponential steps (in the data dimension). Subsequently, we generalize this result to mixture distributions of sub-Gaussian modes. This generalization extends our earlier result on Gaussian mixtures to a significantly larger family of mixture models, as the sub-Gaussian family includes any distribution over an $\ell_{2}$-norm-bounded support set.
(a) Traditional mixture model
(b) Mixture model with in-between mode $P^{(0)}$
Figure 1: Traditional mixture models studied in the literature vs. our analyzed mixture distribution possessing the in-between mode $P^{(0)}$. $P^{(0)}$ is supposed to contain a minor probability mass, yet with a significantly higher variance than the other modes $P^{(1)},\ldots,P^{(k)}$.
@@
To reduce Langevin dynamics’ large iteration complexity shown under a high-dimensional input vector, we propose Chained Langevin Dynamics (Chained-LD). Since Langevin dynamics could suffer from the curse of dimensionality, we decompose the sample ${\mathbf{x}}\in{\mathbb{R}}^{d}$ into $d/Q$ patches ${\mathbf{x}}^{(1)},\cdots,{\mathbf{x}}^{(d/Q)}$, each of constant size $Q$, and sequentially generate every patch ${\mathbf{x}}^{(q)}$ for all $q\in[d/Q]$ statistically conditioned on previous patches, i.e., $P({\mathbf{x}}^{(q)}\mid{\mathbf{x}}^{(0)},\cdots{\mathbf{x}}^{(q-1)})$. The combination of all patches generated from the conditional distribution faithfully follows the probability density $P({\mathbf{x}})$ due to the chain rule, while drawing samples from each patch requires less cost due to the reduced dimension. We also provide a theoretical analysis of Chained-LD by reducing the convergence of a $d$-dimensional sample to the convergence of each patch.
@@
Finally, we present the results of several numerical experiments to validate our theoretical findings. For synthetic experiments, we consider moderately high-dimensional Gaussian mixture models, where the vanilla Langevin dynamics could not find all the components within a million steps, while Chained-LD could capture all the components with correct frequencies in ${\mathcal{O}}(10^{4})$ steps. For experiments on real image datasets, we consider a mixture of two modes by using the original images from MNIST/Fashion-MNIST training dataset (black background and white digits/objects) as the first mode and constructing the second mode by i.i.d. flipping the images (white background and black digits/objects) with probability 0.5. Following from [4], we trained a Noise Conditional Score Network (NCSN) to estimate the score function. Our numerical results indicate that Chained-LD was capable of finding both modes regardless of initialization. We summarize the contributions of this work as follows:
@@
  • Theoretically studying the mode-seeking properties of vanilla Langevin dynamics,
@@
  • Numerically comparing the mode-seeking properties of vanilla and chained Langevin dynamics.
@@ 2 Related Works
Langevin Dynamics: The convergence guarantees for Langevin diffusion, a continuous version of Langevin dynamics, are classical results extensively studied in the literature [12, 13, 14, 15]. Langevin dynamics, also known as Langevin Monte Carlo, is a discretization of Langevin diffusion typically modeled as a Markov Chain Monte Carlo [16]. For unimodal distributions, e.g., the probability density function that is log-concave or satisfies log-Sobolev inequality, the convergence of Langevin dynamics is provably fast [17, 18, 19]. However, for multimodal distributions, the non-asymptotic convergence analysis is much more challenging [20]. [2] gave an upper bound on the convergence time of Langevin dynamics for arbitrary non-log-concave distributions with certain regularity assumptions, which, however, could be exponentially large without imposing more restrictive assumptions. [3] studied the special case of a mixture of Gaussians of equal variance and provided an analysis of sampling from general non-log-concave distributions.
@@
Score-based Generative Modeling: A central task in unsupervised learning involves learning the underlying probability distribution of training data and efficiently generating new samples from the distribution. Since [25] proposed sliced score matching which can train deep models to learn the score functions of implicit probability distributions on high-dimensional data, score-based generative modeling (SGM) has been going through a spurt of growth. Annealed Langevin dynamics [4] estimates the noise score of the probability density perturbed by Gaussian noise and utilizes Langevin dynamics to generate samples from a sequence of decreasing noise scales. [26] conducted an analysis of the effect of noise levels on the performance of annealed Langevin dynamics. Denoising diffusion probabilistic model (DDPM) [27] incorporates a step-by-step introduction of random noise into data, followed by learning to reverse this diffusion process in order to generate desired data samples from the noise. [28] unified anneal Langevin dynamics and DDPM via a stochastic differential equation. A recent line of work focuses on the non-asymptotic convergence guarantees for SGM with an imperfect score estimation under various assumptions on the data distribution [29, 30, 31, 32, 33, 34, 35]. [36] also investigated the KL convergence guarantees for score-based diffusion models. We highlight that a key difference between SGM and our theoretical analysis is that we assume the sampler has direct access to the true score function, whereas SGM typically focuses on learning the score function from training data.
@@ 3.1 Langevin Dynamics
${\,\mathrm{d}}{\mathbf{x}}_{t}=\nabla_{\mathbf{x}}\log P({\mathbf{x}}_{t}){\,\mathrm{d}}t+\sqrt{2}{\,\mathrm{d}}{\mathbf{w}}_{t},$
where ${\mathbf{w}}_{t}$ is the Wiener process on ${\mathbb{R}}^{d}$. Langevin dynamics, a discretization of the SDE for $T$ iterations, is applied to generate samples. Each iteration of Langevin dynamics is defined as
@@
and running Langevin dynamics on the perturbed data distribution $P_{{\sigma}_{t}}({\mathbf{x}})$ with gradually decreasing noise levels $\left\{{{\sigma}_{t}}\right\}_{t\in[T]}$, i.e.,
@@
where $\delta_{t}$ is the step size and $\bm{\epsilon}_{t}\sim{\mathcal{N}}(\bm{0}_{d},{\bm{I}}_{d})$ is Gaussian noise. When the noise level ${\sigma}$ is vanishingly small, the perturbed distribution is close to the true distribution, i.e., $P_{\sigma}({\mathbf{x}})\approx P({\mathbf{x}})$.
Remark 1.
In our theoretical analysis, we assume the sampler has access to the true score function $\nabla_{\mathbf{x}}\log P_{{\sigma}}({\mathbf{x}})$. In some realistic scenarios such as image datasets, since we do not have direct access to the (perturbed) score function, [4] proposed the Noise Conditional Score Network (NCSN) ${\mathbf{s}}_{\bm{\theta}}({\mathbf{x}},{\sigma})$ to jointly estimate the scores of all perturbed data distributions, i.e., $\forall{\sigma}\in\left\{{{\sigma}_{t}}\right\}_{t\in[T]},\;{\mathbf{s}}_{\bm{\theta}}({\mathbf{x}},{\sigma})\approx\nabla_{\mathbf{x}}\log P_{{\sigma}}({\mathbf{x}})$.
@@ 3.2 Multi-Modal Distributions
We remark that sub-Gaussian distributions include a wide variety of distributions such as Gaussian distributions and any distribution within a bounded $\ell_{2}$-norm distance from the mean ${\bm{\mu}}$.
@@ 4 Theoretical Analysis of the Mode-Seeking Properties of Langevin Dynamics
In this section, we theoretically investigate the mode-seeking properties of vanilla Langevin dynamics. We begin with analyzing Langevin dynamics in Gaussian mixtures, and further generalize the results to sub-Gaussian mixtures. We again highlight that in our theoretical analysis, we assume the sampler has access to the score function $\nabla_{\mathbf{x}}\log P({\mathbf{x}})$ of the underlying distribution $P$.
@@ 4.1 Langevin Dynamics in Gaussian Mixtures
To intuitively understand Assumption 1, we first note that the probability density $p({\mathbf{z}})$ of a Gaussian distribution ${\mathcal{N}}({\bm{\mu}},{\nu}^{2}{\bm{I}}_{d})$ decays exponentially in terms of $\frac{\left\|{{\mathbf{z}}-{\bm{\mu}}}\right\|^{2}}{{\nu}^{2}}$. When a state ${\mathbf{z}}$ is sufficiently far from one mode $P^{(i)}$, the probability density of $P^{(i)}$ is dominated by the high-variance component $P^{(0)}$, which implies that the gradient information from $P^{(i)}$ will be masked by $P^{(0)}$. Hence, the dynamics can only visit the universal mode unless the stochastic noise miraculously leads it to the region of another mode. In addition, it can be verified that $\log\left(\frac{{\nu}_{i}^{2}}{{\nu}_{0}^{2}}\right)-\frac{{\nu}_{i}^{2}}{2{\nu}_{0}^{2}}+\frac{{\nu}_{0}^{2}}{2{\nu}_{i}^{2}}$ is a positive constant for ${\nu}_{i}<{\nu}_{0}$, thus the last requirement of Assumption 1 essentially represents $\left\|{{\bm{\mu}}_{i}-{\bm{\mu}}_{0}}\right\|^{2}\leq{\mathcal{O}}(d)$. We formalize the intuition in Theorem 1 and defer the proof to Appendix A.1.
@@
The constants in the notation $\Omega(d)$ are specified in Equations 6 and 7 in the Appendix. We note that $\left\|{{\mathbf{x}}_{t}-{\bm{\mu}}_{i}}\right\|^{2}>\frac{{\nu}_{0}^{2}+{\nu}_{\max}^{2}}{2}d$ is a strong notion of mode-seekingness, since the density of mode $P^{(i)}={\mathcal{N}}({\bm{\mu}}_{i},{\nu}_{i}^{2}{\bm{I}}_{d})$ concentrates around the $\ell_{2}$-norm ball $\left\{{{\mathbf{z}}:\left\|{{\mathbf{z}}-{\bm{\mu}}_{i}}\right\|^{2}\leq{\nu}_{i}^{2}d}\right\}$. This notion can be translated into a lower bound in terms of other distance measures, e.g., total variation distance in the following corollary, whose proof is deferred to Appendix A.2.
Corollary 1.
Under the same assumptions as in Theorem 1, for all time steps $t\in\{0\}\cup[T]$, the distribution $\hat{P}_{t}$ of the generated sample ${\mathbf{x}}_{t}$ by Langevin dynamics at time $t$ satisfies
$\text{TV}(\hat{P}_{t},P)\geq(1-w_{0})\left(1-\frac{T}{\exp(\Omega(d))}\right).$